OpenStudy (anonymous):

how is each function related to its parent function? graph the function by translating the parent function.

OpenStudy (anonymous):

What's the function?

OpenStudy (anonymous):

y=x+5

OpenStudy (anonymous):

The parent function is y = x So y = x+5 is the graph of y = x just raised (translated) 5 units up.

OpenStudy (anonymous):

can you show me please ?

OpenStudy (anonymous):

Graph y = x. Then graph y = x + 5.